ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 90C15, 90C11 Schultz, Rüdiger, University of Leipzig, Germany Mixed-Integer Linear Programming under Uncertainty Deterministic equivalents to mixed-integer linear programs with random parameters are discussed. Beside analytical properties of these models we present solution methods that combine decomposition techniques with traditional algorithms from integer programming.